if known psi(x)=(2-i)e3ix determine |psi(x)|2


1
Expert's answer
2021-10-02T15:45:07-0400

"\psi(x)=(2-i)e^{3ix},"

"\psi^*(x)=(2+i)e^{-3ix},"

"|\psi(x)|^2=\psi\psi^*=(2-i)e^{3ix}(2+i)e^{-3ix}=5\cdot 1=5."